# Enabling On-Device LLMs Personalization with Smartphone Sensing

### v1

@@ Abstract.
This demo presents a novel end-to-end framework that combines on-device large language models (LLMs) with smartphone sensing technologies to achieve context-aware and personalized services. The framework addresses critical limitations of current personalization solutions via cloud-based LLMs, such as privacy concerns, latency and cost, and limited personal sensor data. To achieve this, we innovatively proposed deploying LLMs on smartphones with multimodal sensor data and customized prompt engineering, ensuring privacy and enhancing personalization performance through context-aware sensing. A case study involving a university student demonstrated the proposed framework’s capability to provide tailored recommendations. In addition, we show that the proposed framework achieves the best trade-off in privacy, performance, latency, cost, battery and energy consumption between on-device and cloud LLMs. Future work aims to integrate more diverse sensor data and conduct large-scale user studies to further refine the personalization. We envision the proposed framework could significantly improve user experiences in various domains such as healthcare, productivity, and entertainment by providing secure, context-aware, and efficient interactions directly on users’ devices.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 05–09, 2024; Melbourne, Australiaisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. Introduction
The advent of large language models (LLMs), such as ChatGPT (OpenAI, 2024a), has revolutionized human-machine interaction and impacted people from all walks of life by leveraging vast amounts of data and sophisticated algorithms to provide users with flexibility and personalized content (Cao et al., 2023). Despite their impressive capabilities in understanding, reasoning, and generation, current LLMs face significant limitations, primarily related to privacy and security, since most contemporary LLMs (e.g., ChatGPT, Claude (Anthropic, 2024b), and Gemini (Google, 2024b)), primarily operate in cloud environments which require users to upload their personal data to the cloud, potentially leading to the leakage of sensitive personal information.
@@
Secondly, there are latency and cost issues to consider. Latency can significantly impact user experience, especially when the network is unstable or facing high request volumes and major outrage (OpenAI, 2024c). In critical situations like healthcare monitoring, where real-time data analysis and responses are essential, such latency is unacceptable as it could potentially compromise patient care and safety. When considering cost, cloud-based LLM services are expensive, preventing their extensive usage. For example, calling APIs from servers can cost around 75 USD per million tokens, as seen with Claude 3 Opus (Anthropic, 2024a).
@@
Lastly, the inability of cloud-based LLMs to adapt to real-time contextual data from users poses a significant challenge to personalization. Existing LLMs require pre-collected datasets, hindering their usage in personalization scenarios that rely heavily on streaming data. When users request tasks beyond their boundaries, generic LLMs often fail to provide accurate results. Given these issues, it is reasonable to consider smartphones as the ideal platform for sensing human activities and delivering personalized services, as LLMs are not yet extensively deployed on smartphones for this purpose.
@@
In this demo, we aim to address aforementioned challenges by leveraging on-device LLMs combined with smartphone sensing technologies to enable personalized and context-aware services. The objective is to develop an end-to-end framework for on-device personalization with personal multimodal information and customized prompt engineering to meet individual user needs. By overcoming privacy concerns, latency and cost, and limited personal sensor data, our approach can provide more secure, context-aware, and efficient personalized services that directly proceed on users’ devices, paving the way for broader applications and improved user experiences in various domains such as healthcare, productivity, and entertainment. To verify the proposed framework, we present a case study exploring a day in the life of a university student and discuss the comparison between on-device and cloud LLMs. To the best of our knowledge, this is the first framework to provide on-device LLMs personalization on smartphones for daily uses.
@@ 2.1. On-Device LLMs
There is a growing trend towards creating smaller models for deployment on edge devices such as smartphones and wearable devices (llama.cpp team, 2023; Lin et al., 2023; Liu et al., 2024; Abdin et al., 2024). On-device LLMs refer to LLMs running locally on devices rather than in the cloud, which mitigates the concerns regarding privacy and latency, as the processing occurs locally without connecting to the Internet. Compared to centralized cloud-based LLMs, which can be plagued by latency and bandwidth issues, on-device LLMs provides faster, more reliable, and more efficient processing, leading to quicker, safer and privacy-preserved decision-making. However, deploying LLMs on edge devices presents challenges due to limited computational resources. Open-source models such as Llama-3-8B (AI@Meta, 2024), Phi-3-mini (Abdin et al., 2024), and Gemma-2B (Gemma-team, 2024) have been introduced and explored for deployment on devices such as PCs and smartphones. However, efforts to deploy lightweight LLMs to edge devices are still at an infant stage and applications based on these models have not been comprehensively studied (Wen et al., 2023; Jaganathan et al., 2024; Chen and Li, 2024; Moniz et al., 2024; Li et al., 2024; Carreira et al., 2023; Qin et al., 2023). In comparison, we are the first framework to provide on-device LLMs personalization on smartphones.
@@ 4.2. On-Device and Cloud LLMs Comparison
Both cloud and on-device LLMs can deliver personalized services in the case study. Compared to cloud LLMs, our work focusing on on-device LLMs prioritizes privacy, as all processing is performed on the smartphone. In terms of performance, cloud LLMs can leverage state-of-the-art models, while on-device LLMs may be inferior in accuracy and have a less up-to-date knowledge base. However, we envision future pocket LLMs will be more lightweight, robust and accurate. Regarding latency, cloud LLMs can experience higher latency due to potential network instability or high request volumes and major outrage (OpenAI, 2024c), affecting real-time performance. In comparison, our framework can provide faster and more reliable responses. In terms of cost, on-device LLMs are free, while cloud LLMs require calling APIs from servers such as ChatGPT-4o (OpenAI, 2024b), Claude 3 Opus (Anthropic, 2024a), Gemini 1.5 Pro (Google, 2024a), which can cost from 5 to 75 USD per million tokens. When considering battery consumption, on-device LLMs directly impact the device’s battery life. Our tests indicated an approximately 3% battery drain in 5 minutes on a Google Pixel 8 Pro during our case study on on-device model inference, highlighting one of the future research factors for on-device LLMs personalization. Finally, when examining energy consumption, cloud LLMs consume considerable energy due to their extensive computational requirements. An LLM consumes 0.1 J per token for every billion parameters (Liu et al., 2024). For example, the cloud-based LLM ChatGPT, with 175B parameters, consumes approximately 17.5 J per token, resulting in substantial energy usage. On-device LLMs, while consuming less energy overall, still contribute to energy usage, albeit to a lesser extent. For instance, a 7B-parameter model consumes 0.7 J per token. Moreover, the substantial energy consumption of cloud LLMs has significant environmental implications, with ongoing energy usage contributing to increased carbon dioxide emissions and exacerbating environmental problems. As such, enabling on-device LLMs could make significant contributions for ecological AI.

### v2

@@ Abstract.
This demo presents a novel end-to-end framework that combines on-device large language models (LLMs) with smartphone sensing technologies to achieve context-aware and personalized services. The framework addresses critical limitations of current personalization solutions via cloud LLMs, such as privacy concerns, latency and cost, and limited personal information. To achieve this, we innovatively proposed deploying LLMs on smartphones with multimodal sensor data through context-aware sensing and customized prompt engineering, ensuring privacy and enhancing personalization performance. A case study involving a university student demonstrated the capability of the framework to provide tailored recommendations. In addition, we show that the framework achieves the best trade-off in privacy, performance, latency, cost, battery and energy consumption between on-device and cloud LLMs. To the best of our knowledge, this is the first framework to provide on-device LLMs personalization with smartphone sensing. Future work will incorporate more diverse sensor data and involve extensive user studies to enhance personalization. Our proposed framework has the potential to substantially improve user experiences across domains including healthcare, productivity, and entertainment.
@@
copyright: rightsretainedjournalyear: 2024doi: 10.1145/3675094.3677545conference: Companion of the 2024 ACM International Joint Conference on Pervasive and Ubiquitous Computing ; October 5–9, 2024; Melbourne, VIC, Australiabooktitle: Companion of the 2024 ACM International Joint Conference on Pervasive and Ubiquitous Computing (UbiComp Companion ’24), October 5–9, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-1058-2/24/10ccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. Introduction
The emergence of large language models (LLMs), exemplified by ChatGPT (OpenAI, 2024a), has revolutionized human-machine interaction and impacted people from all walks of life by leveraging vast amounts of data and sophisticated algorithms to provide users with flexibility and personalized content (Cao et al., 2023). Despite their impressive capabilities in understanding, reasoning, and generation, current LLMs face significant limitations, primarily related to privacy and security, since most contemporary LLMs (e.g., ChatGPT, Claude (Anthropic, 2024b), and Gemini (Google, 2024b)), primarily operate in cloud environments which require users to upload their personal data to the cloud, potentially leading to the leakage of sensitive personal information.
@@
Secondly, there are latency and cost issues to consider. Latency can significantly impact user experience, especially when the network is unstable or facing high request volumes and major outrage (OpenAI, 2024c). In critical situations like healthcare monitoring, where real-time data analysis and responses are essential, such latency is unacceptable as it could potentially compromise patient care and safety. When considering cost, cloud LLM services are expensive, preventing their extensive usage. For example, calling APIs from servers can cost around 75 USD per million tokens, as seen with Claude 3 Opus (Anthropic, 2024a).
@@
Lastly, the inability of cloud LLMs to adapt to real-time contextual data from users poses a significant challenge to personalization. Existing LLMs require pre-collected datasets, hindering their usage in personalization scenarios that rely heavily on streaming data. When users request tasks beyond their boundaries, generic LLMs often fail to provide accurate results. Given these issues, it is reasonable to consider smartphones as the ideal platform for sensing human activities and delivering personalized services, as LLMs are not yet extensively deployed on smartphones for this purpose.
@@
In this demo, we aim to address aforementioned challenges by leveraging on-device LLMs combined with smartphone sensing technologies to enable personalized and context-aware services. The objective is to develop an end-to-end framework for on-device personalization with personal multimodal information and customized prompt engineering to meet individual user needs. By overcoming privacy concerns, latency and cost, and limited personal sensor data, our approach can provide more secure, context-aware, and efficient personalized services that directly proceed on users’ devices, paving the way for broader applications and improved user experiences in various domains such as healthcare, productivity, and entertainment. To verify the proposed framework, we present a case study exploring a day in the life of a university student and discuss the comparison between on-device and cloud LLMs. To the best of our knowledge, this is the first framework to provide on-device LLMs personalization with smartphone sensing.
@@ 2.1. On-Device LLMs
There is a growing trend towards creating smaller models for deployment on edge devices such as smartphones and wearable devices (llama.cpp team, 2023; Lin et al., 2023; Liu et al., 2024; Abdin et al., 2024). On-device LLMs refer to LLMs running locally on devices rather than in the cloud, which mitigates the concerns regarding privacy and latency, as the processing occurs locally without connecting to the Internet. Compared to centralized cloud LLMs, which can be plagued by latency and bandwidth issues, on-device LLMs provides faster, more reliable, and more efficient processing, leading to quicker, safer and privacy-preserved decision-making. However, deploying LLMs on edge devices presents challenges due to limited computational resources. Open-source models such as Llama-3-8B (AI@Meta, 2024), Phi-3-mini (Abdin et al., 2024), and Gemma-2B (Gemma-team, 2024) have been introduced and explored for deployment on devices such as PCs and smartphones. Efforts to deploy lightweight LLMs to edge devices are still at an infant stage and applications based on these models have not been comprehensively studied (Wen et al., 2023; Jaganathan et al., 2024; Chen and Li, 2024; Moniz et al., 2024; Li et al., 2024; Carreira et al., 2023; Qin et al., 2023). In comparison, we are the first framework to provide on-device LLMs personalization on smartphones.
@@ 4.2. On-Device and Cloud LLMs Comparison
Both cloud and on-device LLMs can deliver personalized services in the case study. Compared to cloud LLMs, our work focusing on on-device LLMs prioritizes privacy, as all processing is performed on the smartphone. In terms of performance, cloud LLMs can leverage state-of-the-art models, while on-device LLMs may be inferior in accuracy and have a less up-to-date knowledge base. However, we envision future pocket LLMs will be more lightweight, robust and accurate. Regarding latency, cloud LLMs can experience higher latency due to potential network instability or high request volumes and major outrage (OpenAI, 2024c), affecting real-time performance. In comparison, our framework can provide faster and more reliable responses. In terms of cost, on-device LLMs are free, while cloud LLMs require calling APIs from servers such as ChatGPT-4o (OpenAI, 2024b), Claude 3 Opus (Anthropic, 2024a), Gemini 1.5 Pro (Google, 2024a), which can cost from 5 to 75 USD per million tokens. When considering battery consumption, on-device LLMs directly impact the device’s battery life. Our tests indicated approximately 16.5% RAM usage and 3% battery drain in 5 minutes on a Google Pixel 8 Pro during our case study on on-device model inference, highlighting one of the future research factors for on-device LLMs. Finally, when examining energy consumption, cloud LLMs consume considerable energy due to their extensive computational requirements. An LLM consumes 0.1 J per token for every billion parameters (Liu et al., 2024). For example, the cloud LLM ChatGPT, with 175B parameters, consumes approximately 17.5 J per token, resulting in substantial energy usage. On-device LLMs, while consuming less energy overall, still contribute to energy usage, albeit to a lesser extent. For instance, a 7B-parameter model consumes 0.7 J per token. Moreover, the substantial energy consumption of cloud LLMs has significant environmental implications, with ongoing energy usage contributing to increased carbon dioxide emissions and exacerbating environmental problems. As such, enabling on-device LLMs could make significant contributions for ecological AI.